OpenStudy (anonymous):

Can someone help me with this problem? (x^-2)^2(3xy^5)^4 I'm just confused with which power I raise it to?

OpenStudy (anonymous):

\[(x^{-2})^2(3xy^5)^4\]

OpenStudy (anonymous):

first you need to do this \[x^{m^n}=x^{m*n}\]

OpenStudy (anonymous):

and also \[(xyz)^n=x^ny^nz^n\]

OpenStudy (anonymous):